Sum this series of numbers: 1+4+7+10+13+16+....+8938+8941+8944. How do you find the number of integers and how do you solve? I got 1491 integers [1+(n-1)*3 = 8944 and Sum [n/2*(1+8944)]= 13,336,995.
Answer by stanbon(75887)

Sum this series of numbers: 1+4+7+10+13+16+....+8938+8941+8944. How do you find the number of integers and how do you solve? I got 1491 integers [1+(n-1)*3 = 8944 and Sum [n/2*(1+8944)]= 13,336,995.
----------
a(n)=a(1)+(n-1)d
1+(n-1)3=8944
(n-1)3=8943
n-1=2981
n=2982
The last term is the 2982nd term
Sum= (n/2)(1+8944)
Sum=(2982/2)(8945)
Sum=your answer is correct here
Cheers,
Stan H.